Integrate whole root of (x-alpha)(beta-x) with respect to x

Grade:10

1 Answers

moumi roy
7 years ago
let alpha be a &beta be b
∫(x-a)(b-x)dx=∫{x(b+a)-x2-ab}dx=x2(b+a)/2-x3/3-abx+C
since ∫ of xn gives xn+1/n+1